OpenStudy (anonymous):

x=4y y=4x+3 solve for x and y

OpenStudy (anonymous):

help

OpenStudy (anonymous):

y=4(4y)+3 y=16y+3 -15y=3 15y=-3 y=-3/15 y=-1/5 x=4y x=4*(-1/5) x=-4/5

OpenStudy (anonymous):

y=4(4y)+3 y=16y+3 -15y=3 y=-3/15=-1/5 x=4y x=4(-1/5) x=-4/5

hero (hero):

looks like someone is already helping you

OpenStudy (anonymous):

thank you (: